140625 LML Observations

Isaac arrived at the LML site at 11:10am. No sample had been collected although the doors to the fog collector had opened. The top plate, drip tube assembly, and the strand frames were removed for cleaning. The parts were initially sprayed with 409 then thoroughly rinsed with DI water. The collector body was then sprayed lightly with 409 followed by a thorough rinse with a large volume of DI water. The system was then reassembled for collection during the following day.

Leave a Reply

Your email address will not be published. Required fields are marked *